# On Exact Space-Depth Trade-Offs in Multi-Controlled Toffoli Decomposition

### v2

@@ IV.2 Exact Trade-Off between Toffoli Depth and Clean Ancillae
Given $m$ ancilla qubits to implement an $n$-controlled Toffoli gate, we distribute $m$ as $m_{1}+m_{2}$, where $m_{1}$ ancilla qubits are allocated for Step I, and $m_{2}$ ancilla qubits are reserved for Step III. For $m=3$, we set $m_{1}=2$ and $m_{2}=1$. Furthermore, for $m\geq 4$, we assume $m_{2}=2$ to implement the smaller MCT in Step III using the $2$-clean ancilla technique described in khattar24 .
@@
  • In Step II, information from the remaining control qubits is accumulated into $k(<n)$ qubits using the conditionally cleaned ancilla.
  • In Step III, a $(k+m_{1})$-controlled Toffoli gate is implemented using the remaining $m_{2}$ many ancilla, following the $2$-ancilla technique, and the target qubit is modified.
@@
  • Finally, in Step V, the first $m_{1}$ many Toffoli gates are uncomputed to clean up the ancilla, using logical-AND based uncomputation. This step reduces the overall Toffoli count by $m_{1}$ compared to the circuit presented in khattar24 .
@@
where the Toffoli depth for Step I is $1$, and the Toffoli depth for Step II (and Step IV) are $5$ each. Additionally, in Step III, we need to implement a $6$-controlled Toffoli using $m_{2}=2$ ancilla, which requires an additional Toffoli depth of $3$. Consequently, the Clifford + Toffoli decomposition of the $32$-controlled Toffoli gate, utilizing $(3+2)=5$ ancilla qubits, achieves a total Toffoli depth of $1+2(5)+3=14$. In contrast, the $32$-controlled Toffoli decomposition with $2$ ancilla qubits, as presented in khattar24 , results in a Toffoli depth of $19$.
@@
where the Toffoli depth remains $14$. Thus, it can be observed that beyond a certain point, further increases in the number of ancilla qubits do not necessarily lead to a reduction in the Toffoli depth when using conditionally clean ancilla.
@@
The exact Toffoli depth of Step II or Step IV of an $n$-MCT circuit decomposition, implemented above, using $m=m_{1}+m_{2}$ clean ancilla following the conditionally clean ancilla technique, varies with $n$ as follows.
@@
The number of control qubits, $\sigma$, in Step III of an $n$-MCT circuit decomposition, implemented above using $m=m_{1}+m_{2}$ ancilla qubit, varies with $n$ as follows.
@@
The exact Toffoli depth, $\delta$, of the complete $n$-MCT circuit decomposition using $m=m_{1}+m_{2}$ clean ancilla is given by
@@
It is now understood that the depth of $n$-MCT can be achieved in $\mathcal{O}(\log_{2}n)$, which can be explicitly written $c\log_{2}n$, where $c$ needs to be properly estimated, for exact trade-offs which is the main motivation in this paper. In Proposition 3, we have shown that in the MCT decomposition using two clean ancillae with the technique from khattar24 , $c$ is strictly greater than 3; in fact, it is around 4. However, using the ancilla - Toffoli depth trade-off, $c$ can be further reduced to a certain extent. In the following subsection, we demonstrate that with the conditionally clean ancilla technique, $c$ always remains strictly greater than 1, regardless of the number of ancilla used.
IV.3 Proving the Lower Bound on Toffoli Depth using Conditionally Clean Ancilla
In this subsection, we show that the Toffoli depth of an $n$-MCT using the conditionally clean ancilla technique can never be reduced to $\lceil\log_{2}n\rceil$.
@@ IV.3 Proving the Lower Bound on Toffoli Depth using Conditionally Clean Ancilla
    Assuming that the control states do not need to be restored to their original values, using the conditionally clean ancilla technique, the exact Toffoli depth must be strictly greater than $\lceil\log_{2}n\rceil$, irrespective of the number of available ancilla.
@@
Therefore, Case II holds for all $k\geq 2$ for some $k\in\mathbb{N}$. In conclusion, using the conditionally clean ancilla technique, the exact Toffoli depth of an $n$-controlled Toffoli decomposition can not be reduced to $\lceil\log_{2}n\rceil$, even when the control qubits are not restored to their original states.
@@
The above theorem shows that using the conditionally clean ancilla, the constant factor $c$ can never be reduced to $1$. In the next section, we demonstrate that the exact Toffoli depth in the Clifford + Toffoli decomposition of an $n$-MCT is lower bounded by $\lceil\log_{2}n\rceil$, regardless of the technique or the number of ancilla qubits used. This bound is achievable through complete binary tree decomposition of $n$-controlled Toffoli gates.
@@ VI Conclusion
In this paper, we revisited the $n$-controlled Toffoli decomposition using the conditionally clean ancilla technique described by Khattar and Gidney khattar24 and proposed an exact trade-off between Toffoli depth and the availability of clean ancilla qubits. By leveraging additional ancilla qubits, we achieved a lower Toffoli depth compared to their approach. Furthermore, we demonstrated that the conditionally clean ancilla technique cannot reduce the Toffoli depth strictly to $\lceil\log_{2}n\rceil$, irrespective of unlimited availability of clean ancilla.
@@
Acknowledgments: The authors like to thank Dr. Alessandro Luongo (CQT, NUS, and Inveriant Pte. Ltd., Singapore) for his valuable inputs while preparing this paper. Subhamoy Maitra acknowledges the funding support provided by the “Information Security Education and Awareness (ISEA) Project phase - III, Cluster - Cryptography, initiatives of MeitY, Grant No. F.No. L-14017/1/2022-HRD”.

### v3

@@ IV.2 Exact Trade-Off between Toffoli Depth and Clean Ancillae
Given $m$ ancilla qubits to implement an $n$-controlled Toffoli gate, we distribute $m$ as $m_{1}+m_{2}$, where $m_{1}$ ancilla qubits are allocated for Step I, and $m_{2}$ ancilla qubits are reserved for Step III. For $m=3$, we set $m_{1}=2$ and $m_{2}=1$. Furthermore, for $m\geq 4$, we assume $m_{2}=2$ to implement the smaller MCT in Step III using the $2$-clean ancillae technique described in khattar24 .
@@
  • In Step II, information from the remaining control qubits is accumulated into $k(<n)$ qubits using the conditionally cleaned ancillae.
  • In Step III, a $(k+m_{1})$-controlled Toffoli gate is implemented using the remaining $m_{2}$ many ancillae, following the $2$-ancilla technique, and the target qubit is modified.
@@
  • Finally, in Step V, the first $m_{1}$ many Toffoli gates are uncomputed to clean up the ancillae, using logical-AND based uncomputation. This step reduces the overall Toffoli count by $m_{1}$ compared to the circuit presented in khattar24 .
@@
where the Toffoli depth for Step I is $1$, and the Toffoli depth for Step II (and Step IV) are $5$ each. Additionally, in Step III, we need to implement a $6$-controlled Toffoli using $m_{2}=2$ ancillae, which requires an additional Toffoli depth of $3$. Consequently, the Clifford + Toffoli decomposition of the $32$-controlled Toffoli gate, utilizing $(3+2)=5$ ancilla qubits, achieves a total Toffoli depth of $1+2(5)+3=14$. In contrast, the $32$-controlled Toffoli decomposition with $2$ ancilla qubits, as presented in khattar24 , results in a Toffoli depth of $19$.
@@
where the Toffoli depth remains $14$. Thus, it can be observed that beyond a certain point, further increases in the number of ancilla qubits do not necessarily lead to a reduction in the Toffoli depth when using conditionally clean ancillae.
@@
The exact Toffoli depth of Step II or Step IV of an $n$-MCT circuit decomposition, implemented above, using $m=m_{1}+m_{2}$ clean ancillae following the conditionally clean ancillae technique, varies with $n$ as follows.
@@
The number of control qubits, $\sigma$, in Step III of an $n$-MCT circuit decomposition, implemented above using $m=m_{1}+m_{2}$ ancilla qubits, varies with $n$ as follows.
@@
The exact Toffoli depth, $\delta$, of the complete $n$-MCT circuit decomposition using $m=m_{1}+m_{2}$ clean ancillae is given by
@@
It is now understood that the depth of $n$-MCT can be achieved in $\mathcal{O}(\log_{2}n)$, which can be explicitly written $c\log_{2}n$, where $c$ needs to be properly estimated, for exact trade-offs which is the main motivation in this paper. In Proposition 3, we have shown that in the MCT decomposition using two clean ancillae with the technique from khattar24 , $c$ is strictly greater than 3; in fact, it is around 4. However, using the ancilla - Toffoli depth trade-off, $c$ can be further reduced to a certain extent. In the following subsection, we demonstrate that with the conditionally clean ancillae technique, $c$ always remains strictly greater than 1, regardless of the number of ancilla qubits used.
IV.3 Proving the Lower Bound on Toffoli Depth using Conditionally Clean Ancillae
In this subsection, we show that the Toffoli depth of an $n$-MCT using the conditionally clean ancillae technique can never be reduced to $\lceil\log_{2}n\rceil$.
@@ IV.3 Proving the Lower Bound on Toffoli Depth using Conditionally Clean Ancillae
    Assuming that the control states do not need to be restored to their original values, using the conditionally clean ancillae technique, the exact Toffoli depth must be strictly greater than $\lceil\log_{2}n\rceil$, irrespective of the number of available ancilla.
@@
Therefore, Case II holds for all $k\geq 2$ for some $k\in\mathbb{N}$. In conclusion, using the conditionally clean ancillae technique, the exact Toffoli depth of an $n$-controlled Toffoli decomposition can not be reduced to $\lceil\log_{2}n\rceil$, even when the control qubits are not restored to their original states.
@@
The above theorem shows that using the conditionally clean ancillae, the constant factor $c$ can never be reduced to $1$. In the next section, we demonstrate that the exact Toffoli depth in the Clifford + Toffoli decomposition of an $n$-MCT is lower bounded by $\lceil\log_{2}n\rceil$, regardless of the technique or the number of ancilla qubits used. This bound is achievable through complete binary tree decomposition of $n$-controlled Toffoli gates.
@@ VI Conclusion
In this paper, we revisited the $n$-controlled Toffoli decomposition using the conditionally clean ancilla technique described by Khattar and Gidney khattar24 and proposed an exact trade-off between Toffoli depth and the availability of clean ancilla qubits. By leveraging additional ancilla qubits, we achieved a lower Toffoli depth compared to their approach. Furthermore, we demonstrated that the conditionally clean ancillae technique cannot reduce the Toffoli depth strictly to $\lceil\log_{2}n\rceil$, irrespective of unlimited availability of clean ancilla.
@@
Acknowledgments: The authors like to thank Dr. Alessandro Luongo (CQT, NUS, and Inveriant Pte. Ltd., Singapore) for his valuable inputs while preparing this paper. Suman Dutta and Anupam Chattopadhyay acknowledge the support of ‘MoE AcRF Tier 1 award RT10/23’. Subhamoy Maitra acknowledges the funding support provided by the “Information Security Education and Awareness (ISEA) Project phase - III, Cluster - Cryptography, initiatives of MeitY, Grant No. F.No. L-14017/1/2022-HRD”.